OpenStudy (anonymous):

log xy^2/Z is equalivent to 2logxy-logz, is this true? if not, expand it

jimthompson5910 (jim_thompson5910):

I'm going to use these log rules http://www.purplemath.com/modules/logrules.htm \[\Large \log\left(\frac{xy^2}{z}\right)\] \[\Large \log\left(xy^2\right)-\log\left(z\right) ... \text{Rule 2}\] \[\Large \log\left(x\right)+\log\left(y^2\right)-\log\left(z\right) ... \text{Rule 1}\] \[\Large \log\left(x\right)+2\log\left(y\right)-\log\left(z\right) ... \text{Rule 3}\] ------------------------------------------------------------------------- So \[\Large \log\left(\frac{xy^2}{z}\right)\] is equivalent to \[\Large \log\left(x\right)+2\log\left(y\right)-\log\left(z\right)\]

OpenStudy (anonymous):

thanks!

jimthompson5910 (jim_thompson5910):

yw